the mass of a typical car is m= 1000 kg, so its weight on earths surface is mg=9800N. suppose you have two containers, one with N electrons and another with N protons. these two containers are placed a distance 10m apart, and the electric force between them is equal to the weight of the car. Find N


1
Expert's answer
2022-05-14T16:44:44-0400

"F=\frac{N^2ke^2}{r^2},"

"N=\frac re\sqrt\frac Fk=6.5\cdot 10^{16}."